Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 26
Entry time: 01 Apr 2019, 10:21
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Report after dispatch meeting 09.30 |
|
|
SPB/SFX
- The quadrant motion system test for the AGIPD presented some issues with the limit switches. Being fixed.
- Julabo chiller: waiting for feedback from Julabo to know whether the chiller is going to be repaired on time for next users. If not, SPB/SFX are trying to borrow one from CFEL (necessary parts availability to be checked.)
- CSS loop restarted
- CRL chamber will be brought down tomorrow. Cables in the rackroom will be re-shuffled by tomorrow evening.
FXE
SQS
- NQS chamber is being cabled up
- Crates for Loop 8 and 11 being installed by AE
- VAC interlock definitions are being updated.
- Solution for pnCCD has been found. To control it an existing PLC and cables will be used. Draft of the proposed solution will be sento to AE by TB.
SCS
- Loop 13 firmware export will be provided to AE by EET on friday.
MID
- Preparation of proposal for next users is ongoing (including AGIPD and EPIX).
- Cabling issues for EPIX are being solved. Calibration of the detector will follow
- One crate for epix has been changed and is being mounted back today
- OPT VAC is now remote and being tested
- Tomorrow CY will organize a review meeting for the epix where also needed people for the tests foreseen for 4-5.04 will be addressed.
- Tech. commissioning of the CHILLER2 (AGIPD) will be done tomorrow
HED
- Cabling phase II vs loop operation: All the PRIO cables have been promised by the compoany to be installed by next friday.
- Early next week LW will make a statement whether Loop 5 can be put in operation already during the week.
- Installation of imagers behind the mono - Karen is organizing it. VAC supports.
OPT Lasers
XRO
- Cooling installation in the SA1/3 mirrors ig going well. Commissioning foreseen for next WE. SQS and SCS shutters would need to be opened. Will be discussed offline with the instruments.
- Minor works ongoing.
XPD
VAC
- Hirex 2 has been installed
- Support ongoing
- Today MID mono will be opened.
AE
- SASE3 and 1 tunnel MOV loops have been updated. Feedback to AE is needed!
- Issue with dymaic triggers has been fixed.
ITDM
- XFEL gateway will be updated in 15 minutes
- Jddd usage examples will be posted by ITDM
DET
EET
CAS
- Tunnels updated to Karabo 2.4.0. For the instruments coordination with scientist is ongoing (SCS already done)
AoB
- Mail sent to the GLs asking for feedback concerning shutdown activities by GW.
|